OpenStudy (anonymous):

The radius of the Earth is approximately 6,370 kilometers. What is the distance in scientific notation? 637.0 × 10^1 kilometers 6.370 × 10^3 kilometers 63.70 × 10^2 kilometers 637.0 × 10^2 kilometers will medal and fan if correct

OpenStudy (whpalmer4):

\[10^1= 10\]\[10^2=10*10=100\]\[10^3=10*10*10=1000\] \[637.0*10^1=637*10=\]\[6.37*10^3=6.37*1000=\]\[63.70*10^2=63.70*100=\]\[637.0*10^2=637.0*100=\] Several of those give you 6370, so which one is correct? Remember that scientific notation has the mantissa (the part before the \(*10^n\)) within a range such that the magnitude of the mantissa is 1 or greater, but also less than 10.